If the kinetic energy of a body changes by 20% then its momentum would change by –
(A) 20% (B) 24% (C) 40% (D) 44%

Grade:

2 Answers

Arun
3 years ago
Momentum is proportional to speed, so when momentum is increased by 20%, or multiplied by a factor 1.2 , then then speed must have increased by 20%. Kinetic energy is proportional to the square of speed, so with speed increasing by a factor 1.2 , Kinetic Energy wil increase by the square of the factor 1.2, which is 1.44 or 44%
Vikas TU
3 years ago

We know that

 
  • K=(1/2)*m*v^2
  • p=m*v

    so you can see the relation by computing this two.

    Let us massage equation 2 a bit, for convenience:

     
    • p=m*v v=p/m

      Now we can plug this into our Kinetic energy equation, by replacing v for p/m, which yields:

       
      • K=(1/2)*m*(p/m)^2

        which, after manipulating a little will shorten to:

         
        • K=(1/2m)*p^2

          And now we have a relation between K and p. Let us now see what happens when you increase p by 20%.

          An increase by 20% in p can be expressed as 1.2*p . So if you replace, in the last equation, p by 1.2*p, you will end up with

           
          • K=(1.44/2m)*p^2

            So, if you increase the momentum by 20%(1.2), the Kinetic energy will increase by a factor of 1.44 or, in percentage, by 44%.

            Kind regards,